# Advancing Prompt Learning through an External Layer

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.isbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3680953ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
@@ 1. Introduction
Vision-language models (VLMs), such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), have demonstrated remarkable generalization performance for various downstream tasks. VLMs are typically trained to align textual and visual modalities using large-scale datasets, which allows them to encode open-vocabulary (Fang et al., 2021, 2020) concepts in a shared embedding space. Thanks to the modality matching ability, CLIP has achieved remarkable success across various downstream tasks (Greer et al., 2024; Luo et al., 2018, 2023c; Zhou et al., 2022a; Luo et al., 2023d; Guo et al., 2024), such as action recognition (Wang et al., [n. d.]), image generation (Sanghi et al., 2021), and image segmentation (Luo et al., 2023a; Zheng Ding, 2023). One of the attractive features of CLIP is the ability to perform zero-shot inference, where some pre-defined text inputs (a.k.a. prompts) are used to generate classification weights for predicting image features during inference.
Seminal explorations involve hand-crafted text prompts, e.g., “a photo of a [class]” as the prompt for text encoder. Advanced works introducing a set of learnable parameters to adapt VLMs to downstream tasks. For instance, CoOp (Zhou et al., 2022c) keeps the weights of CLIP frozen while learning the text embeddings of prompts for efficient task-specific adaptation. These prompt-learning approaches achieve significant performance improvements over manually tuned prompts on various typical scenarios (Chen et al., 2023b; Cao et al., 2024c; Etchegaray et al., 2024). However, the learned text embeddings typically produce worse performance compared with hand-crafted prompts (Radford et al., 2021) for unseen tasks. Specifically, applying these learned text prompts to unseen classes leads to degenerated model generalization performance. Recent works mitigate this problem through the overfitting view. To overcome the challenge, an image-conditional prompt (CoCoOp) learning approach (Zhou et al., 2022b) is proposed to improve the generalization performance on unseen classes. Unfortunately, CoCoOp exhibits lower generalization performance than hand-crafted zero-shot CLIP when applied to novel classes, as depicted in Table 4. This could be attributed to the learned embeddings are usually invalid on novel classes.
@@ 2.1. Preliminaries
Figure 2. Comparison of our method with CoOp and MaPLe. Our method freezes the textual embeddings and learn an External Layer (EnLa) on top of it. Optimal transport is introduced to align and visual and textual modalities apart from the conventional cross entropy (CE) loss. The textual and visual encoders are connected through a strengthening feature. This feature and the learnable visual prompt are fused and utilized for deep prompting of the visual encoder.
@@
where ⓒ is the concatenating operation. Follow CoOp (Zhou et al., 2022c), the embeddings $\mathbf{w}$ can be shared across classes. In the inference phase, the prompts with the learned embeddings can produce textual features for classification.
@@ 2.2. EnLa with Frozen Text Embeddings
Frozen Text Embeddings. Built upon the observation, we propose to freeze the text embeddings of prompts for unseen tasks. Specifically, the text embeddings are set to be frozen and non-learnable, which aims to release the potential generalization ability of CLIP and ensure pre-trained ability for unseen tasks. Here, the input words in the prompt are typically initialized using ‘a photo of a’, which can be vectorized into embeddings. In our upcoming ablation experiments, we plan to incorporate other manual template inputs for validation. However, freezing text embeddings of prompts will cause limited model capacity, which results in a lack of capacity to adapt VLMs for downstream tasks.
@@ 2.3. Alignment with Optimal Transport
Specifically, let $\boldsymbol{v}$ denote the visual feature and $\boldsymbol{t}$ denote the text feature. The output of the image encoder is a tensor with the shape, where $H$ and $W$ are the height and width. Therefore, we can obtain $M=H\times W$ local visual features. Further, let $\boldsymbol{V}=\left\{\left.\boldsymbol{v}_{m}\right|_{m=1}^{M}\right\}$ denote the local visual features as the fixed set, let $\boldsymbol{t}_{k}$ denote the text feature as the fixed set for class ${k}$. Our method learns the transport plan $\mathbf{T}$ by minimizing the following OT distance to push $\boldsymbol{t}_{k}$ to $\boldsymbol{V}$ for fine-grained alignment:
@@ 2.4. Alignment with Strengthening Feature
We introduce learnable visual embeddings for enhancing learnability of VLMs for downstream tasks. In contrast to VPT (Jia et al., 2022), our method employ a multi-modal prompt design. Specifically, let $\mathcal{T}_{\epsilon}(\cdot)$ denote the V-L function parameterized by $\epsilon$ to transfer features from textual branch to visual branch. The $\mathcal{T}_{\epsilon}(\cdot)$ is a function net with dimensionality of [512, 768] for dimension alignment of the visual branch and text branch. The input feature of the V-L function is the output produced by the EnLa. Further, we append $L$ learnable visual input embeddings $\mathcal{P}_{v}$, which is a vector with [$L$, 768], given as follows,
@@
To improve generalization capabilities, we are conducting an evaluation to assess the performance of different connection positions of the image encoder, as depicted in Table 2.
@@
The hidden layer of the image encoder (row-1) combines the benefits of prompting in both branches by enforcing the image encoder representation ability through text knowledge transfer to the deep layer of the image encoder without visual embeddings initialization layer, and the input layer of the image encoder (row-2) is based on the connection of the visual embeddings initialization. The position (row-2) is more effective than the position (row-1) in unseen tasks. It can be attributed to that position (row-1) lacks visual embedding initialization, leading to a limited ability to adapt the image encoder with the unbalance of V-L alignment. As a consequence, in our subsequent comprehensive experiments, we adopted the connection position with the input layer of the image encoder (row-2) in our approach. Let $\boldsymbol{e}$ denote the output of EnLa, and let $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})$ denote the fusion embeddings, it can be expressed as:
(13) $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})=\boldsymbol{p}_{\boldsymbol{v}}+\mathcal{T}_{\boldsymbol{\epsilon}}(\boldsymbol{e}).$
The fusion embeddings are further introduced in the transformer of the image encoder for deep prompting.
@@ 3.1. Benchmark Setting
We compare the performance with CLIP (Radford et al., 2021), CoOp (Zhou et al., 2022c), Clip-Adapter (Gao et al., 2023), CoCoOp (Zhou et al., 2022b), PLOT (Chen et al., 2022a), and MaPLe (Khattak et al., 2023). Please note that CLIP was originally a zero-shot method. However, in this context, we linearly process it using a few-shot method. The Clip-Adapter (Gao et al., 2023) is a frozen prompt method with the adapter component behind the encoder, which is different from EnPrompt. We compared it with PLOT (Chen et al., 2022a), which is an optimal transport method. PLOT optimizes four sentences simultaneously (“a photo of a dog”, “a picture of a dog”, “a drawing of a dog”, “a good drawing of a dog”). The output text feature of EnPrompt is global feature, the input sentence is one text prompt such as “a photo of a”, and the text prompt is non-learnable.
@@
Implementation Details. We use a ViT-B/16-based CLIP model, and report results averaged over 3 runs. We set the learnable visual embedding length to 4. We use deep prompting with an image encoder. Training for 50 epochs for few-shot setting, 20 epochs for domain generalization setting and cross-dataset evaluation setting. In the base-to-novel setting, we use 30 epochs for most datasets. We use an SGD optimizer with a learning rate of 0.0025. All experiments are run on a single Nvidia A6000 GPU. For domain generalization and cross-dataset evaluation, we train the ImageNet source model on all classes with K = 16 shots in the first 3 transformer layers of the image encoder. For the few-shot learning and base-to-novel tasks, we set visual embeddings learning depth to 9. We initialize the text embeddings using the hand-crafted words of ‘a photo of a’.
@@ 4.1. Comparison of Different Templates
‘a drawing of a’ 98.3 77.10 86.42 $\pm$0.05
‘a painting of the’ 97.2 77.7 86.39 $\pm$0.05
‘a photo of a’ 98.47 77 86.43 /
@@ 4.2. Training Strategy
In Table 8, we evaluate the performance of different EnLa training strategies as an ablation. In our approach, the EnLa is implemented as a neural network, which is essential to learn all potentially valuable features during the training stage in order to achieve effective generalization (Zhang et al., 2021b; Fang et al., 2024; Li et al., 2021; Zhang et al., 2020b; Muandet et al., 2013; Chen et al., 2022b, 2023a; Zhang et al., 2024). To this end, we focus on training the model for more epochs to learn richer features (row-2), resulting in improved generalization performance.
@@
EnLa (2 epoch) 83.22 75.91 79.68
EnLa (30 epoch) 84.71 76.90 80.64
@@ 4.5. Analyzing of Multi-modal Methods
To validate the effect of EnPrompt (Ours), we conducted the HM task and few-shot learning task using two-modal designs and the method completely opposite to EnPrompt in Table 9. We observed that the method (row-2) has a lower performance in novel classes than the method (row-3), which is due to the learnable text embeddings of method (row-2) being weak in generalization compared to the frozen text prompt. Finally, we observed that the method (row-3) and method (row-2) performed better in the novel classes than method (row-1) while still maintaining a non-generalization few-shot learning task. This is attributed to the fact that method (row-1) is to have neither learnable visual embeddings nor learnable text embeddings. It indicates that prompt learning is crucial for adapting pre-trained VLMs for generalization.
@@ 4.6. Inference Stage Computational Cost
In Table 10, we show the compute cost analysis of EnPrompt (Ours) and compare it with text embedding learning approaches. Although EnPrompt uses the EnLa, its overall parameters exceed only by $0.52\%$ over CLIP. Compared to MaPLe, EnPrompt has fewer parameters and lower coupling. In terms of inference speed, CoCoOp is significantly slower. In contrast, EnPrompt has no such overhead, obtaining a higher performance with less training time. Although CoOp has a small number of parameters, due to the mismatch of V-L alignment, the training time of 10 epochs for CoOp is similar to ours. EnPrompt is more simpler than text multi-prompt input method (PLOT) in textual design component. Further, EnPrompt provides better convergence as it gets better HM as compared to MaPLe in 10 epochs.
@@ 5.1. Prompt Learning in Vision Language Models
Recently, the strong generalization capability of CLIP (Radford et al., 2021) has made it a foundation for many methods (Fang et al., 2019; Cao et al., 2024b, a) in adapting pre-trained VLMs for downstream tasks (Ding et al., 2023; Luo et al., 2023b; Li et al., 2023b, a). Prompt learning (Liu et al., 2021; Lester et al., 2021; Li and Liang, 2021) is a widely used technique in NLP for learning downstream tasks. The use of text prompts, which are instructions provided to the language branch of a Vision-Language model (VLM), is a common practice to enhance task understanding (Gao et al., 2023; Zhang et al., 2021a; Yang et al., 2024b). CoOp (Zhou et al., 2022c) and CoCoOp (Zhou et al., 2022b) fine-tune the CLIP model specifically for few-shot image recognition by optimizing a continuous set of token embeddings in the language branch. The image-conditional prompt of CoCoOp significantly contributes to enhancing generalization to unseen classes (Fang et al., 2023, 2022a, 2022b) and mitigating the risk of overfitting to the limited labeled data. Moreover, some approaches (Yao et al., 2023a; Lu et al., 2022; Zhu et al., 2023; Yao et al., 2023b) constrain the proposed learnable prompts to contain the essential general knowledge and prior distribution learning. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022b) ensures that the language model attends to relevant visual information when generating predictions. In addition to single-modal prompt tuning (Jia et al., 2022), some approaches (Khattak et al., 2023; Liu et al., 2023) introduce the multi-modal prompt-tuning designs in CLIP to effectively align V-L representations (Pan et al., 2024; Luo et al., 2024; Di et al., 2024). However, their methods have not fully released the potential generalization ability of CLIP.
@@ 5.2. Optimal Transport
Optimal Transport (OT) has recently gained significant attention in various theoretical and application tasks to measure the distance between two probability distributions over metric spaces (Guo et al., 2022; Arjovsky et al., 2017; Tanwisuth et al., 2021; Balaji et al., 2020). For instance, Redko et al.(Redko et al., 2019) tackle the target shift problem by aligning domain distributions using the OT framework. PLOT (Chen et al., 2022a) introduce a local text feature with textual multi-prompt input. Ours obtain a global text feature for textual one prompt, instead of a local text feature with textual multi-prompt input.

### v6

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680953isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
@@ 1. Introduction
Vision-language models (VLMs), such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021), have demonstrated remarkable generalization performance for various downstream tasks (Xie et al., 2024, 2023, 2022). VLMs are typically trained to align textual and visual modalities using large-scale datasets, which allows them to encode open-vocabulary (Fang et al., 2021, 2020) concepts in a shared embedding space. Thanks to the modality matching ability, CLIP has achieved remarkable success across various downstream tasks (Greer et al., 2024; Luo et al., 2018, 2023c; Zhou et al., 2022a; Luo et al., 2023d; Guo et al., 2024), such as action recognition (Wang et al., [n. d.]), image generation (Sanghi et al., 2021), and image segmentation (Luo et al., 2023a; Zheng Ding, 2023). One of the attractive features of CLIP is the ability to perform zero-shot inference, where some pre-defined text inputs (a.k.a. prompts) are used to generate classification weights for predicting image features during inference.
Seminal explorations involve hand-crafted text prompts, e.g., “a photo of a [class]” as the prompt for text encoder. Advanced works introduce a set of learnable parameters to adapt VLMs to downstream tasks. For instance, CoOp (Zhou et al., 2022c) keeps the weights of CLIP frozen while learning the text embeddings of prompts for efficient task-specific adaptation. These prompt-learning approaches achieve significant performance improvements over manually tuned prompts on various typical scenarios (Chen et al., 2023; Cao et al., 2024c; Etchegaray et al., 2024). However, the learned text embeddings typically produce worse performance compared with hand-crafted prompts (Radford et al., 2021) for unseen tasks. Specifically, applying these learned text prompts to unseen classes leads to degenerated model generalization performance. Recent works mitigate this problem through the overfitting view. To overcome the challenge, an image-conditional prompt (CoCoOp) learning approach (Zhou et al., 2022b) is proposed to improve the generalization performance on unseen classes. Unfortunately, CoCoOp exhibits lower generalization performance than hand-crafted zero-shot CLIP when applied to novel classes, as depicted in Table 4. This could be attributed to the fact that learned embeddings are usually invalid on novel classes.
@@ 2.1. Preliminaries
Figure 2. Comparison of our method with CoOp and MaPLe. Our method freezes the textual embeddings and learns an External Layer (EnLa) on top of it. Optimal transport is introduced to align visual and textual modalities apart from the conventional cross entropy (CE) loss. The textual and visual encoders are connected through a strengthening feature. This feature and the learnable visual prompt are fused and utilized for deep prompting of the visual encoder.
@@
where ⓒ is the concatenating operation. Following CoOp (Zhou et al., 2022c), the embeddings $\mathbf{w}$ can be shared across classes. In the inference phase, the prompts with the learned embeddings can produce textual features for classification.
@@ 2.2. EnLa with Frozen Text Embeddings
Frozen Text Embeddings. Built upon the observation, we propose to freeze the text embeddings of prompts for unseen tasks. Specifically, the text embeddings are set to be frozen and non-learnable, which aims to release the potential generalization ability of CLIP and ensure pre-trained ability for unseen tasks. Here, the input words in the prompt are typically initialized using “a photo of a”, which can be vectorized into embeddings. In our upcoming ablation experiments, we plan to incorporate other manual template inputs for validation. However, freezing text embeddings of prompts will cause limited model capacity, which results in a lack of capacity to adapt VLMs for downstream tasks.
@@ 2.3. Alignment with Optimal Transport
Specifically, let $\boldsymbol{v}$ denote the visual feature and $\boldsymbol{t}$ denote the textual feature. The output of the image encoder is a tensor with the shape, where $H$ and $W$ are the height and width. Therefore, we can obtain $M=H\times W$ local visual features. Further, let $\boldsymbol{V}=\left\{\left.\boldsymbol{v}_{m}\right|_{m=1}^{M}\right\}$ denote the local visual features as the fixed set, let $\boldsymbol{t}_{k}$ denote the text feature as the fixed set for class ${k}$. Our method learns the transport plan $\mathbf{T}$ by minimizing the following OT distance to push $\boldsymbol{t}_{k}$ to $\boldsymbol{V}$ for fine-grained alignment:
@@ 2.4. Alignment with Strengthening Feature
We introduce learnable visual embeddings for enhancing learnability of VLMs for downstream tasks. In contrast to VPT (Jia et al., 2022), our method employs a multi-modal prompt design. Specifically, let $\mathcal{T}_{\epsilon}(\cdot)$ denote the V-L function parameterized by $\epsilon$ to transfer features from textual branch to visual branch. The $\mathcal{T}_{\epsilon}(\cdot)$ is a function net with dimensionality of [512, 768] for dimension alignment of the visual branch and text branch. The input feature of the V-L function is the output produced by the EnLa. Further, we append $L$ learnable visual input embeddings $\mathcal{P}_{v}$, which is a vector with [$L$, 768], given as follows,
@@
To analyze generalization capabilities of alignment, we are conducting an evaluation to assess the performance of different connection positions of the image encoder, as depicted in Table 2.
@@
The deep layer (row-1) of the image encoder is the hidden layer of the image encoder, without visual input (first) layer. The visual input layer (row-2) of the image encoder method is focused on the initialization of the visual embeddings. These designs combine prompting benefits in both branches by enforcing the image encoder representation ability through text knowledge transformation. The position (row-2) is more effective than the position (row-1) in unseen tasks. It can be attributed to that position (row-1) lacks visual embedding initialization, leading to a limited ability to adapt the image encoder with the unbalance of V-L alignment. As a consequence, in our subsequent comprehensive experiments, we adopted the connection position with the input layer of the image encoder (row-2) in our approach. Let $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})$ denote the fusion embeddings, it can be expressed as:
(13) $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})=\boldsymbol{p}_{\boldsymbol{v}}+\mathcal{T}_{\boldsymbol{\epsilon}}(\boldsymbol{e}),$
the $\boldsymbol{e}$ is the output of EnLa, and the fusion embeddings are further introduced in the transformer of the image encoder for deep prompting.
@@ 3.1. Benchmark Setting
We compare the performance with CLIP (Radford et al., 2021), CoOp (Zhou et al., 2022c), Clip-Adapter (Gao et al., 2023), CoCoOp (Zhou et al., 2022b), PLOT (Chen et al., 2022), and MaPLe (Khattak et al., 2023). Please note that CLIP was originally a zero-shot method. However, in this context, we linearly process it using a few-shot method. The Clip-Adapter (Gao et al., 2023) is a frozen prompt method with the adapter component behind the encoder, which is different from EnPrompt. We compared it with PLOT (Chen et al., 2022), which is an optimal transport method. PLOT optimizes four sentences simultaneously (“a photo of a dog”, “a picture of a dog”, “a drawing of a dog”, “a good drawing of a dog”). The output text feature of EnPrompt is global feature, the input sentence is one text prompt such as “a photo of a”, and the text prompt is non-learnable.
@@
Implementation Details. We use a ViT-B/16-based CLIP model, and report results averaged over 3 runs. We set the learnable visual embedding length to 4. We use deep prompting (Khattak et al., 2023) for the image encoder. Training for 50 epochs for few-shot setting, 20 epochs for domain generalization setting and cross-dataset evaluation setting. In the base-to-novel setting, we apply 30 epochs. We use an SGD optimizer with a learning rate of 0.0025. All experiments are run on a single Nvidia A6000 GPU. For domain generalization and cross-dataset evaluation, we train the ImageNet source model on all classes with K = 16 shots in the first 3 transformer layers of the image encoder. For the few-shot learning and base-to-novel tasks, we set visual embeddings learning depth to 9. We initialize the text embeddings using the hand-crafted words of “a photo of a []”.
@@ 4.1. Comparison of Different Templates
“a drawing of a” 98.3 77.10 86.42 $\pm$0.05
“a painting of the” 97.2 77.7 86.39 $\pm$0.05
“a photo of a” 98.47 77 86.43 /
@@ 4.2. Training Strategy
In Table 8, we evaluate the performance of different EnLa training strategies as an ablation. In our approach, the EnLa is implemented as a neural network, which is essential to learn all potentially valuable features during the training stage in order to achieve effective generalization (Li et al., 2021; Zhang et al., 2020b, 2024). To this end, we focus on training the model for more epochs to learn richer features (row-2), resulting in improved generalization performance (Zhang et al., 2021b; Fang et al., 2024).
@@
1: EnLa (2 epoch) 83.22 75.91 79.68
2: EnLa (30 epoch) 84.71 76.90 80.64
@@ 4.5. Analyzing of Multi-modal Methods
To validate the effect of EnPrompt (Ours), we conducted the base-to-novel task and non-generalization few-shot task using two-modal designs (row-1) and the method completely opposite to EnPrompt (row-2) in Table 9. The method (row-2) is a visual external layer based on frozen visual embeddings with learning textual embeddings. We observed that the method (row-2) has a lower performance in novel classes than the method (row-3), which is due to the learnable text embeddings of method (row-2) being weak in pre-trained generalization compared to the frozen text prompt. Finally, we observed that the method (row-3) and method (row-2) performed better in the novel classes than method (row-1) while still maintaining a non-generalization few-shot task. This is attributed to the fact that method (row-1) has neither learnable visual embeddings nor learnable text embeddings. It indicates that prompt learning is crucial for adapting pre-trained VLMs for generalization.
@@ 4.6. Inference Stage Computational Cost
In Table 10, we show the compute cost analysis of EnPrompt (Ours) and compare it with text embedding learning approaches. Although EnPrompt uses the EnLa, its overall parameters exceed only by $0.52\%$ over CLIP. Compared to MaPLe, EnPrompt has fewer parameters and lower coupling. In terms of inference speed, CoCoOp is significantly slower. In contrast, EnPrompt has no such overhead, obtaining a higher performance with less training time. Although CoOp has a small number of parameters, due to the mismatch of V-L alignment, the training time of 10 epochs for CoOp is similar to ours. EnPrompt is simpler than textual multi-prompt input method (PLOT) in textual design component. Further, EnPrompt provides better convergence as it gets better HM as compared to MaPLe in 10 epochs.
@@ 5.1. Prompt Learning in Vision Language Models
Recently, the strong generalization capability of CLIP (Radford et al., 2021) has made it a foundation for many methods (Fang et al., 2019; Cao et al., 2024b, a) in adapting pre-trained VLMs for downstream tasks (Ding et al., 2023; Luo et al., 2023b; Li et al., 2023b, a). Prompt learning (Liu et al., 2021; Lester et al., 2021; Li and Liang, 2021) is a widely used technique in NLP for learning downstream tasks. The use of text prompts, which are instructions provided to the language branch of a Vision-Language model (VLM), is a common practice to enhance task understanding (Gao et al., 2023; Zhang et al., 2021a; Yang et al., 2024b). CoOp (Zhou et al., 2022c) and CoCoOp (Zhou et al., 2022b) fine-tuning the CLIP model specifically for few-shot image recognition by optimizing a continuous set of token embeddings in the language branch. The image-conditional prompt of CoCoOp significantly contributes to enhancing generalization to unseen classes (Fang et al., 2023, 2022a, 2022b) and mitigating the risk of overfitting to the limited labeled data. Moreover, some approaches (Yao et al., 2023a; Lu et al., 2022; Zhu et al., 2023; Yao et al., 2023b) constrain the proposed learnable prompts to contain the essential general knowledge and prior distribution learning. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022b) ensures that the language model attends to relevant visual information when generating predictions. In addition to single-modal prompt tuning (Jia et al., 2022), some approaches (Khattak et al., 2023; Liu et al., 2023) introduce the multi-modal prompt-tuning designs in CLIP to effectively align V-L representations (Pan et al., 2024; Luo et al., 2024; Di et al., 2024). However, their methods have not fully released the potential generalization ability of CLIP.
@@ 5.2. Optimal Transport
Optimal Transport (OT) has recently gained significant attention in various theoretical and application tasks to measure the distance between two probability distributions over metric spaces (Guo et al., 2022; Arjovsky et al., 2017; Tanwisuth et al., 2021; Balaji et al., 2020). For instance, Redko et al.(Redko et al., 2019) tackle the target shift problem by aligning domain distributions using the OT framework. PLOT (Chen et al., 2022) introduces a local text feature with textual multi-prompt input. Ours obtain a global text feature for textual one prompt, instead of a local text feature with textual multi-prompt input.